OpenStudy (anonymous):

Lana makes $8 an hour plus $12 an hour for every hour of overtime. Overtime hours are any hours more than 40 hours for the week. Part A: Create an equation that shows the amount of wages earned, W, for working x hours in a week when there is no overtime. (3 points) Part B: Create an equation that shows the amount of wages earned, S, for working y hours of overtime. Hint: Remember to include in the equation the amount earned from working 40 hours. (3 points) Part C: Lana earned $368 in 1 week. How many hours (regular plus overtime) did she work? (4 points)

OpenStudy (anonymous):

please help me!

OpenStudy (studio-games):

Ok. Lemme see here.

OpenStudy (anonymous):

ok

OpenStudy (studio-games):

This FLVS?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

can you please help me out?

OpenStudy (studio-games):

Oh ok neat. I do it too. Lemme see if I have this answered in my questions alwready.

OpenStudy (anonymous):

ok

OpenStudy (studio-games):

Ok so I dont seee it so here. Part A: She works for x hrs (we are ignoring the given hours for now) a week, right? She also earns $8 an hour. You are looking for w, so it's going to be w = what?

OpenStudy (anonymous):

8

OpenStudy (studio-games):

Good

OpenStudy (studio-games):

Part B: So, we are looking for how much she earns (now represented by "S," don't let that confuse you, we are just changing the variable) PLUS hours working overtime (represented by y). It says to not forget how much she earns for working her regular hours. To find that amount, solve the FIRST equation we made for 40 hrs. Just plug in 40 where "x" is and find the answer.

OpenStudy (anonymous):

i don't know how to

OpenStudy (anonymous):

wait what equation the x=8?

OpenStudy (studio-games):

Ok let me see. So we need to know what she earns

OpenStudy (studio-games):

Represented by S

OpenStudy (anonymous):

ok

OpenStudy (studio-games):

Plus the hours she is working.

OpenStudy (studio-games):

In overtime.

OpenStudy (anonymous):

yeah

OpenStudy (studio-games):

It says to not forget how much she earns for working her regular hours. To find that amount, solve the FIRST equation we made for 40 hrs. Just plug in 40 where "x" is and find the answer.

OpenStudy (anonymous):

so what would be the equation then

OpenStudy (anonymous):

?

OpenStudy (anonymous):

can you just give me the answer with an explanation

OpenStudy (studio-games):

So lets work together on it then.

OpenStudy (anonymous):

ok

OpenStudy (studio-games):

Ok give me a second.

OpenStudy (studio-games):

Internet is VERY slow for some reason.

OpenStudy (anonymous):

k

OpenStudy (studio-games):

w=8(40) = 320 Now for S = ?

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

just tell me what part a part b and part c are then we discuss it please

OpenStudy (studio-games):

Ok

OpenStudy (studio-games):

let me write it down then

OpenStudy (anonymous):

ok

OpenStudy (studio-games):

Part A: W=8x where x equals amount of hours, and W equals wages. Part B: S=8x+12(y) where y equals overtime hours and S equals wages Part C: 368=8(40)+12(y) -320=12(y) 48=12y 48/12=y 4=y x=40, y=4. X+Y=44 hours

OpenStudy (studio-games):

There enjoy.

OpenStudy (anonymous):

ok is there anyway we can communicate like by kick or something?

OpenStudy (anonymous):

and i have one last question

OpenStudy (anonymous):

Solve: negative 1 over 2 x + 3 = –x + 7

OpenStudy (studio-games):

I have Skype and google hangouts.

OpenStudy (studio-games):

Ok solving it.

OpenStudy (studio-games):

MInd medaling and fanning me please?

OpenStudy (studio-games):

X = 8

OpenStudy (studio-games):

Enjoy! Have a good day!

OpenStudy (anonymous):

yes i will medal you thanks

OpenStudy (studio-games):

Thanks!

OpenStudy (anonymous):

whats your Skype?